OpenStudy (anonymous):

Find the vertices of the ellipse x^{2} +4y^{2} = 1

OpenStudy (mertsj):

\[\frac{x^2}{1}+\frac{y^2}{\frac{1}{4}}=1\]

OpenStudy (mertsj):

\[a=1, b = \frac{1}{2}\]

OpenStudy (anonymous):

ok thx thats all i need i just wasn't sure if that was right :D

OpenStudy (mertsj):

yw